The average annual snowfall for Boise, ID, is 19.8 inches. Anchorage, AK, receives an average of 74.5 inches.
1. Anchorage receives, on average, ___% more snow than Boise.
2. Boise receives, on average, ___% less snow than Anchorage.

Answer by Theo(13342)

the average for boise, idaho is 19.8 inches.

the average for anchorage, alaska is 74.5 inches.

the absolute value of the difference is 74.5 - 19.8 = 54.7 inches.

54.7 / 19.8 = 2.762626263 * 100 = 276.2626263 percent.

54.7 / 74.5 = .7342281879 * 100 = 73.42281879 percent.

the average snowfall in anchorage, alaska is 276.26% greater than the average snowfall in boise, idaho.

the average snowfall in boise, idaho is 73.42% less than the average snowfall in anchorage alaska.

the solutions were rounded to 2 decimal digits.

you should round as you are required to do.